A positive charge of 6.6 µC and a negative charge of 4.0 µC exert a force of -25.0 N on each other. How far apart are the charges?
1
Expert's answer
2020-02-24T10:47:11-0500

"\vec{F_{21}}=-\vec{F_{12}}"

"|\vec{F_{21}}|=|\vec{F_{12}}|=F"

Behind Coulomb's Law

"F=k \cdot \frac{|q_1|\cdot |q_2|}{r^2}"

Where will we write

"r=\sqrt{ \frac{k \cdot |q_1|\cdot |q_2|}{F}}"

"r=\sqrt{ \frac{9 \cdot10^9 \cdot 6.6 \cdot 10^{-6}\cdot 4.0 \cdot 10^{-6}}{25}}=0.097 m=9.7 sm"